# Preprocessing Disks for Convex Hulls, Revisited

### v1

@@ 3.1 Preliminaries
Supersequences and sorting.
A universal word is a string that contains as subsequences all permutations of its distinct characters. The question of constructing the shortest universal word was first posed by Knuth and attributed to Karp [9]. The length of known shortest universal words are in [1]. The best known upper bound is due to Tan [37]. An earlier bound of $n^{2}-7/3n+19/3$ was given by Radomirovic [31]. His proof is constructive and also gives an $O(n^{2})$ algorithm to construct such a word. Uznański shows that testing whether a given string is a universal word is coNP-complete [38].
The problem of finding a sorted sequence in a supersequence is closely related to the longest increasing subsequence problem. Fredman shows that the problem has a $\Omega(n\log n)$ lower bound [18]. There is also a matching $O(n\log n)$ dynamic programming algorithm, which is commonly taught in advanced algorithms courses.
3.2 Preprocessing algorithm
We will use the following result.
Lemma 3.6 ([32]).
The length of the shortest sequence on $m$ symbols which contains all permutations of $[1,m]$ as a subsequence is upper-bounded by $m^{2}$. Moreover, such a sequence of length $m^{2}$ can be computed in $O(m^{2})$ time.
While Radomirocić [32] does not discuss the running time, it is implicit in their construction. Ultimately we will apply it when $m$ is the ply, which we assume is bounded. Using this result, we prove the following technical lemma, which we will also reuse later in Section 2.
Lemma 3.7.
Let $\cal I$ be a set of $n$ unit intervals in $\mathbb{R}$ of ply $\Delta$. There is a $(3,2\Delta)$-smooth sorting-supersequence $\Xi$ of $\cal I$ with at most $4\Delta n$ intervals.
Figure 5: An example of a sequence of unit intervals of ply $\Delta=3$.
Note, to get a sequence of length $n\Delta^{2}$, we could simply apply Lemma 3.6 to every cell of the subdivision induced by $I$. However, Lemma 3.7 gives a more refined bound.
Proof 3.8.
Assume the intervals =${\cal I}=\{I_{1},\ldots,I_{n}\}$ are ordered and indexed from left to right (since they are unit intervals, this order is well-defined). Now, let $S_{1}=\{I_{i}\mid I_{i}\cap I_{1}\neq\emptyset\}$ be the set of all intervals that intersect the leftmost interval (including the leftmost interval itself) and let $\eta_{1}=\min\{i\mid I_{i}\cap I_{1}=\emptyset\}$ be the index of the first interval not in $S_{1}$. We recursively define $S_{i+1}$ to be the set of all intervals that intersect $I_{\eta_{i}}$ and are not in $S_{i}$. Then, let $W_{i}=\bigcup_{I\in S_{i}}I$ be the window from the left endpoint of the leftmost interval in $S_{i}$ to the right endpoint of the rightmost interval in $S_{i}$.
Now, define $S^{\prime}_{i}$ to be the set of all intervals from $S_{i}$ unioned with all intervals from $S_{i+1}$ which intersect $W_{i}$. We claim that $|S^{\prime}_{i}|\leq 2|S_{i}|-1\leq 2\Delta-1$. This is because any interval from $S_{i+1}$ intersecting $W_{i}$, that is not already in $S_{i}$, must contain the right endpoint of $W_{i}$, and as $\Delta$ is the ply, there can therefore be at most $\Delta-1$ such intervals. We further claim that every pair of intervals that intersect each other is present in some set $S^{\prime}_{i}$: indeed, an interval from $S_{i}$ cannot intersect an interval from $S_{i+2}$ since then the interval in $S_{i+2}$ would also intersect the leftmost interval of $S_{i+1}$ (but then it would be part of $S_{i+1}$).
Let $k$ be the resulting number of $S_{i}^{\prime}$ sets. We process these sets of intervals from left to right, and apply Lemma 3.6 to $S^{\prime}_{i}$ for $i=1,\ldots,k$. Let the resulting $i$th sequence be $\Xi_{i}$. Then we simply concatenate the resulting sequences, to get a sequence $\Xi=\Xi_{1}\cup\Xi_{2}\cup\cdots\cup\Xi_{k}$.
As argued above, $|S_{i}^{\prime}|\leq 2\Delta$ and moreover $\sum_{i=1}^{k}|S_{i}^{\prime}|\leq 2\sum_{i=1}^{k}|S_{i}|=2n$. By Lemma 3.6, $|\Xi_{i}|=|S_{i}^{\prime}|^{2}$. Thus $|\Xi|$ is bounded by a sum of squares of values, i.e. the $|S_{i}^{\prime}|$, which are each bounded by $2\Delta$ and sum to at most $2n$. It is not hard to see that such a sum of squares is maximized when $S_{i}^{\prime}=2\Delta$ for all $i$ and $k=n/\Delta$. Thus the length of $\Xi$ is at most $n/\Delta\cdot 4\Delta^{2}=4n\Delta$.
Now, let $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim{\cal I}$ be any realization, listed in sorted left to right order on the real line. We divide the points in $P$ into $k$ groups: let $P_{i}\subset P$ be the set of all points that are contained in $W_{i}$ but not in $W_{i-1}$. Let $\Pi=\Pi(P)$ be the ordered string of intervals corresponding to the ordered points of $P$, and let $\Pi_{i}$ be the part of $\Pi$ corresponding to $P_{i}$. The string $\Pi_{i}$ is a substring of $\Xi_{i}$ (by Lemma 3.6). Therefore, the points in $P_{i}$ can be charged to the subsequence $\Xi_{i}$, thus proving $\Xi$ is a sorting-supersequence, as the $\Xi_{i}$ are concatenated in order.
Finally, we need to argue that the resulting sequence $\Xi$ is $(3,2\Delta)$-smooth. To argue that $\alpha=3$, note that all intervals in $S_{i}$ intersect each other, and thus all intervals in $S^{\prime}_{i}$ either intersect each other or intersect a common interval. In the first case, the distance between any two points in them is at most $2$; in the latter case they could have points at distance $3$ from each other. Moreover, the $S_{i}^{\prime}$ are sorted on the real line (say by the left ends of their leftmost interval), and since we concatenate the $\Xi_{i}$ in this order, this bound for $\alpha$ on each $\Xi_{i}$ holds for $\Xi$ as well. To argue that $\beta=2\Delta$, note that the largest number of unit intervals that can intersect an interval of length $J$ is $(|J|+1)\cdot\Delta$, where $(|J|+1)\cdot\Delta\leq 2\Delta|J|$ when $|J|>1$.
We are now ready to finish the preprocessing algorithm.
Proof 3.9 (Proof of Theorem 3.3).
Observe that Lemma 3.7 establishes the existence of the desired sorting-supersequence. Moreover, the proof of Lemma 3.7 is constructive. First, sort the $n$ unit intervals in $O(n\log n)$ time, yielding the $S_{i}$ sets. $S_{i}^{\prime}$ can then be computed (using this sorted ordering) from $S_{i}$ in $O(\Delta)$ time, and thus $O(\Delta n)$ time over all $i$. Finally, Lemma 3.6 states that the time to compute $\Xi_{i}$ from $S_{i}^{\prime}$ is proportional to $|\Xi_{i}|$ (which itself is $|S_{i}^{\prime}|^{2}$). Thus $\Xi$ is computed from the $S_{i}^{\prime}$ in $O(\Delta n)$ time overall, as it is the concatenation of the $\Xi_{i}$, and Lemma 3.7 already argued $|\Xi|\leq 4\Delta n$.
4 Reconstruction in One Dimension
We now consider the problem of recovering the sorted order of a set of values from a supersequence as computed in Section 3. We will first consider the problem in the classical preprocessing framework, where the goal is to achieve linear-time reconstruction, in Section 4.1, and show that for smooth sequences this is achievable in the Real RAM, while on the Word RAM we do not require the sequences to be smooth. Then, in Section 4.2, we discuss how to combine the idea of supersequences with the aim of achieving sublinear reconstruction.
4.1 Standard Reconstruction
For the reconstruction phase, we are now given a sorting-supersequence $\Xi$ of $\mathcal{R}$, and additionally a set of points (values) $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$, and the goal is to efficiently compute the sorted order of $P$ using $\Xi$. (In particular, in linear time when $\Xi$ is $(O(1),O(1))$-smooth.)
We will assume that the regions in $\Xi$ come as pointers to the original regions in $\mathcal{R}$, and that our sets $\mathcal{R}$ and $P$ are equipped with bidirectional pointers.
The natural approach to the reconstruction problem is to replace each region in $\Xi$ by its corresponding point in $P$. This results now in a sequence $X$ of numbers, which comes with a guarantee that the sorted order is a subsequence of $X$.
Definition 4.1.
Let $X=x_{1},\ldots,x_{n}$ be a sequence of $n$ numbers (possibly with duplicate numbers), which contains $k\leq n$ unique numbers. Then we call $X$ a hidden sorted sequence if it contains a subsequence of length $k$ with all $k$ distinct numbers in sorted order.
First, we show that given such a sequence $X$, we can find the sorted subsequence in linear time when universal hashing is allowed. Subsequently, we show even without universal hashing linear time is possible when the sequence is appropriately smooth. It is an open problem whether linear time is possible without universal hashing nor a smoothness bound.
Sorting hidden sorted sequences in linear time in the word RAM model.
We first show that there is a natural greedy algorithm to sort hidden sorted sequences, which runs in expected linear time assuming universal hashing.
Lemma 4.2.
Given a hidden sorted sequence $X=x_{1},\ldots,x_{n}$ with $k$ unique numbers, there is a greedy algorithm which outputs a sorted subsequence of length $k$ containing all $k$ distinct numbers. This algorithm runs in $O(n)$ expected time if universal hashing is allowed, and otherwise it runs in $O(n\log k)$ time.
Proof 4.3.
Create a stack which initially contains just $x_{1}$. Now process the items in the sequence in order (starting at $x_{2}$). Let $x_{i}$ be the current item being processed. If $x_{i}$ exists in the stack already then discard it. Otherwise, pop items off the stack until the top item is less than $x_{i}$, at which point we push $x_{i}$ onto the stack. [^4] Note that the above algorithm is oblivious of the value $k$.
Let $x_{i_{1}},\ldots,x_{i_{k}}$ be any subsequence of $X$ containing all $k$ unique items in sorted order. To argue correctness, we prove by induction that after $x_{i_{j}}$ has been processed by our algorithm, the bottom of the stack contains the $j$ smallest (unique) items in sorted order. (For $j<k$ there may be additional items above on the stack, but for $j=k$ this is not possible as by construction the stack contains sorted unique items.) For the base case when $j=1$, the claim trivially holds as then the smallest item has been processed (either when $x_{i_{j}}$ is processed or earlier), at which point the algorithm will place it at the bottom of the stack, and the algorithm will never remove it. Now consider some $j>1$. By induction, after $x_{i_{j-1}}$ was processed, the stack contains the $j-1$ smallest (unique) items in sorted order at the bottom, and these can never be removed by the algorithm. Thus as $x_{i_{j}}$ is the next smallest item, the same logic as in the base case implies that after $x_{i_{j}}$ is processed, it will be on the stack immediately after the first $j-1$ items, and will never be removed.
Assuming we store the current stack items in a hash table, it takes expected constant time to check if $x_{i}$ is already in the stack. Thus we get a linear time algorithm, since in each time step we are either inserting or deleting from the stack, and each item is inserted at most once. If hashing is not allowed, we can instead use a binary tree for the items in the stack (whose size is at most $k$), and thus the running time is $O(n\log k)$.
Note that the algorithm runs in linear time when hashing is available, but as we are interested in working in the Real RAM model, hashing is not available to us. We note that the problem of finding the sorted subsequence in a hidden sorted sequence is a special case of the longest increasing subseqeunce problem, which has a $\Omega(n\log n)$ lower bound in the decision tree model; however, in our problem if their are no duplicates the sequence is already sorted, whereas this is not the case for LIS, and thus it is unclear whether a lower bound for LIS implies a lowerbound for our problem. Whether a hidden sorted sequence can be recovered in linear time in the Real RAM model is left as an open problem.
Sorting smooth sequences
For our purposes, it is sufficient to recover the sorted subsequence from a hidden sorted sequence that additionally is $(\alpha,\beta$)-smooth. We assume $\alpha$ and $\beta$ are constants, so the goal is to sort the sequence in time of the form $O(f(\alpha,\beta)\cdot n)$.
Note that if we convert the supersequence $\Xi$ as constructed in Section 3 to their corresponding values in $P$, we get a sequence that is $(3,2\Delta)$-smooth.
Lemma 4.4.
Given a $(\alpha,\beta)$-smooth hidden sorted sequence $X=x_{1},\ldots,x_{n}$ with $k$ unique numbers we can output a sorted subsequence of length $k$ containing all $k$ distinct numbers in $O(n\log(\alpha\beta))$ time in the Real RAM model.
Proof 4.5.
Recall the algorithm of Lemma 4.2, which greedily processes the items in $X$ in order, discarding $x_{i}$ if it has already occurred, and otherwise popping items off the stack until the top item is less than $x_{i}$, at which point we push $x_{i}$ onto the stack. We already argued this algorithm is correct, though its running time depends on how long it takes to determine whether $x_{i}$ has already occurred.
The claim is that if $x_{i}$ has already occurred, then is must be one of the largest $\alpha\beta$ items seen so far. Assuming this claim is true, we can implement this algorithm in $O(n\log(\alpha\beta))$ time, as maintaining a balanced binary tree with the largest $\alpha\beta$ items seen, takes $O(\log(\alpha\beta))$ time per round (i.e. $O(n\log(\alpha\beta))$ overall), and the remaining stack operations are linear time overall.
To prove the claim, fix any $x_{i}$, and let $x_{j}$ be the largest value item such that $j<i$. If $x_{i}>x_{j}$, the $x_{i}$ has not been seen before, and the claim trivially holds. So suppose that $x_{i}\leq x_{j}$. Since $X$ is $(\alpha,\beta)$-smooth, we have that $x_{j}-x_{i}\leq\alpha$. Thus if $x_{i}$ has occurred already, it must be within an interval of length $\alpha$ (ending at $x_{j}$), and there can be at most $\alpha\beta$ distinct points in such an interval, by the packing property of $(\alpha,\beta)$-smooth sequences, thus implying the claim.
Proof 4.6 (Proof of Theorem 3.4).
If $\Xi$ is an $(\alpha,\beta)$-smooth sorting-supersequence of $\mathcal{R}$, then by definition for any $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$, if we replace the regions from $\Xi$ by their corresponding points in $P$, then it results in an $(\alpha,\beta)$-smooth hidden sorted sequence. Thus Lemma 4.4, directly states we can recover the sorted order of $P$ in $O(|\Xi|\log(\alpha\beta))$ time.
4.2 Sublinear reconstruction
The preprocessing framework was originally developed with linear-time reconstruction as the ultimate goal: since, during the reconstruction phase, as we need to spend linear time to replace each region in $\mathcal{R}$ by the corresponding point in $P$, we cannot hope to be faster. Van der Hoog et al. [39, 40] argue that this viewpoint may be limiting, and that in some applications we may be content with an output structure that still contains some regions, as long as we are guaranteed that the true points in those regions are combinatorially in the correct location. They argue that the preprocessing framework may be adapted to allow for this behaviour, at the cost of introducing another phase: the reconstruction phase is formally separated into a first subphase (that can take sublinear time), in which the auxiliary structure $\Xi$ is transformed into another structure $\Xi^{\prime}$ which is combinatorially equivalent to the desired output $S(P)$, and a second subphase in which $\Xi^{\prime}$ is actually transformed into $S(P)$ in linear time, if so desired.
We show that the same behavior can be quite naturally obtained when the auxiliary structure $\Xi$ is a supersequence.
Definition 4.7.
A marked sorting-supersequence $\Xi$ of $\mathcal{R}$ is a sequence of (possibly recurring) intervals of $\mathcal{R}$, some of which may be marked, and such that, for any point set $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$, the sorted order of $P$ is a subsequence of $\Xi$, that contains all marked items of $\Xi$.
A normal sorting-supersequence is a marked supersequence with no items marked. To achieve maximum benefit, we want to mark as many items as possible, since these may be skipped during reconstruction.
In the case of sorting, intervals are marked if the position in the sorted order does not depend on the location of the point in the interval. This happens exactly when the interval is disjoint from all other intervals.
We adapt the preprocessing algorithm as follows. In $O(n\log n)$ time, we sort the intervals (by their left endpoint) as before. Now, in linear time, we scan through the intervals and detect any intervals that do not overlap with any other interval. We mark these intervals, and collect clusters of remaining intervals that are separated by the marked intervals. We apply Lemma 3.7 to each cluster separately, and concatenate the resulting sequences.
In order to be able to quickly skip over marked intervals during the reconstruction phase, we can equip the resulting sorting-supersequence with pointers from each item to the next non-marked item. During the reconstruction phase we then also apply Lemma 4.4 separately to each subsequence of consecutive non-marked items, in time propertional to the total number of non-marked items in the list. This results in a mixed sequence of sorted points and marked intervals, which is guaranteed to be in the correct sorted order (this list would be called $\Xi^{\prime}$ in [39]).
5 Preprocessing Disjoint Unit Disks for the convex hull
Our goal in this section is to prove Theorems 2.10 and 2.11 for the case of disjoint unit disks.
5.1 Definitions and notation
Throughout this section, let =$\mathcal{R}=\{R_{1},\ldots,R_{n}\}$ be a set of $n$ disjoint unit disks in the plane, arbitrarily indexed. Let ()=$\mathcal{B}(\mathcal{R})=\{B_{1},\ldots,B_{k}\}$ be the subset of disks from $\mathcal{R}$ on the boundary of the convex hull of $\mathcal{R}$, indexed in counterclockwise order of appearance on the convex hull. We refer to $\mathcal{B}(\mathcal{R})$ as the boundary disks, and disks in the set $\mathcal{I}(\mathcal{R})=\mathcal{R}\setminus\mathcal{B}(\mathcal{R})$ as interior disks. Refer to Figure 6. When $\mathcal{R}$ is clear from the context, we may write simply $\mathcal{B}$ and $\mathcal{I}$ instead of $\mathcal{B}(\mathcal{R})$ and $\mathcal{I}(\mathcal{R})$.
Figure 6: The convex hull of =$\mathcal{R}=\{R_{1},\ldots,R_{11}\}$, with yellow and blue disks on the boundary (depending on whether potential or guaranteed, see Figure 2), orange internal disks, and strips consisting of two consecutive boundary disks. In this case ()=$\mathcal{B}(\mathcal{R})=\{B_{1},\ldots,B_{6}\}=\{R_{7},R_{2},R_{5},R_{9},R_{11},R_{4}\}$.
For any $i\in[1,k]$, define the $i$th strip as $s_{i}=\mathrm{CH}(B_{i},B_{i+1})$, where $\mathrm{CH}$ denotes the convex hull (and $s_{k}=\mathrm{CH}(B_{k},B_{1})$). We refer to the chain of strips, ordered by increasing $i$, as the strip hull of $\mathcal{R}$. (Indeed the convex hull of all strips is equal to that of all regions, where straight segments on the boundary follow individual strips and turns occur at the intersection of the common disk with the next strip.) Furthermore, for any $i$, we call the line segment connecting the centers of $B_{i}$ and $B_{i+1}$ the spine of the corresponding strip. Observe that the spines themselves define an inner convex polygon, which we call the spine hull.
@@ 5.2 Geometric preliminaries
Proof 5.2 (Proof of Lemma 5.1).
@@
Proof 5.4 (Proof of Corollary 5.3).
@@
Proof 5.6 (Proof of Corollary 5.5).
Consider any point $z$ in the plane. $z$ is contained in a strip a strip if and only if a unit disk $D$ centered at $z$ intersects the spine of that strip. So consider the convex polygon $P$ which is the spine hull of the strips. Observe that the edges of $P$ have length at least 2, as the defining disks of a strip are disjoint. Thus applying Lemma 5.1 with $x=2$, we know that $D$ intersects at most 4 edges of $P$, and therefore $z$ intersects at most 4 strips. The ply of the arrangement is the maximum of all points in the plane, and since $z$ was an arbitrary point, the ply is at most 4.
@@ 5.2.1 Fréchet and Hausdorff
We use the following definition of Fréchet distance from [25, Definition 3.10], which applies both to the unit interval $\mathbb{S}^{0}$, and well as the unit circle $\mathbb{S}^{1}$.
@@
$\delta_{F}(P,Q):=\inf_{\alpha:\mathbb{S}^{k}\to\mathbb{S}^{k},\beta:\mathbb{S}^{k}\to\mathbb{S}^{k}}\max_{t\in\mathbb{S}^{k}}||P(\alpha(t))-Q(\beta(t))||,$
@@
[25, Theorem 5.1] discusses the equivalence between Hausdorff distance and Fréchet distance for closed convex curves, originally argued in [4].
@@
5.3 Proof of Theorem 2.10 (Existence) for Disjoint Disks
In this section, we argue that a sequence $\Xi$, as described in Theorem 2.10, exists. We first argue that we can collect the disks that intersect a single strip.
@@
Proof 5.15 (Proof of Lemma 5.14).
@@
    every vertex of the $Q$-hull is in distance 1 from the spine hull, or, in other words, every vertex of the $Q$-hull is contained in one of the strips, which follows from Observation 6.
@@
We will now prove Theorem 2.10 for the case of disjoint unit disks; that is, there exists a $(3\sqrt{2},12)$-smooth quarter-hull-supersequence $\Xi$ for $\mathcal{R}$. In fact, the sequence in the case of disjoint disks is even $(1,3)$-smooth; we will argue the precise smoothness constants only for the general bounded ply case in Section 6.
Proof 5.16 (Proof of Theorem 2.10 for disjoint disks).
@@
5.4 Proof of Theorem 2.11 (Preprocessing) for Disjoint Disks
In this section, we argue that we can also compute such a sequence $\Xi$ in $O(n\log n)$ time. First, we want to argue that we can compute the arrangement of all disks and strips in $O(n\log n)$ time.

### v2

@@ 4.1 Preliminaries
We begin by discussing additional definitions and preliminaries for the preprocessing phase.
4.1.1 Definitions and Notation
Throughout this section, let $\mathcal{R}=\{R_{1},\ldots,R_{n}\}$ be a set of $n$ disjoint unit disks in the plane, arbitrarily indexed. We recall some definitions from Section 2. Let $\mathcal{B}(\mathcal{R})=\{B_{1},\ldots,B_{k}\}$ be the subset of boundary disks (Definition 2.3), indexed in counterclockwise order of appearance on the convex hull, see Figure 7. When $\mathcal{R}$ is clear from the context, we may write simply $\mathcal{B}$.
Figure 7: The convex hull of $\mathcal{R}=\{R_{1},\ldots,R_{11}\}$, with yellow and blue disks on the boundary (depending on whether potential or guaranteed, see Figure 2), orange internal disks, and strips consisting of two consecutive boundary disks. In this case $\mathcal{B}(\mathcal{R})=\{B_{1},\ldots,B_{6}\}=\{R_{7},R_{2},R_{5},R_{9},R_{11},R_{4}\}$.
For any $i\in[1,k]$, define the $i$th strip as $s_{i}=\mathrm{CH}(B_{i},B_{i+1})$ (where $s_{k}=\mathrm{CH}(B_{k},B_{1})$). We refer to the chain of strips, ordered by increasing $i$, as the strip hull of $\mathcal{R}$. Furthermore, for any $i$, we call the line segment connecting the centers of $B_{i}$ and $B_{i+1}$ the spine of the corresponding strip. Observe that the spines themselves define an inner convex polygon [^5] A stronger requirement than was the case for potentially non-unit disks in Section 2., which we call the spine hull.
@@ 4.1.2 Geometric Properties
Proof 4.2 (Proof of Lemma 4.1).
@@
Proof 4.4 (Proof of Corollary 4.3).
@@
Proof 4.6 (Proof of Corollary 4.5).
Consider any point $z$ in the plane. $z$ is contained in a strip if and only if a unit disk $D$ centered at $z$ intersects the spine of that strip. So consider the convex polygon $P$ which is the spine hull of the strips. Observe that the edges of $P$ have length at least 2, as the defining disks of a strip are disjoint. Thus applying Lemma 4.1 with $x=2$, we know that $D$ intersects at most 4 edges of $P$, and therefore $z$ intersects at most 4 strips. The ply of the arrangement is the maximum of all points in the plane, and since $z$ was an arbitrary point, the ply is at most 4.
@@ 4.1.3 Fréchet and Hausdorff
We use the following definition of Fréchet distance from [27, Definition 3.10], which applies both to the unit interval $\mathbb{S}^{0}$, and well as the unit circle $\mathbb{S}^{1}$.
@@
[…]
@@
[27, Theorem 5.1] discusses the equivalence between Hausdorff distance and Fréchet distance for closed convex curves, originally argued in [2].
@@
4.2 Preprocessing Disjoint Unit Disks
Our goal in this section is to prove Theorems 3.4 for the case of disjoint unit disks.
4.2.1 Existence of a Supersequence for Disjoint Unit Disks
In this section, we argue that a sequence $\Xi$, as described in Theorem 3.4, exists.
Theorem 4.12.
Let $\mathcal{R}$ be a set of $n$ disjoint unit disks in the plane. There exists a $(1,3)$-smooth quarter-hull-supersequence $\Xi$ of $\mathcal{R}$ of size $O(n)$.
We first argue that we can collect the disks that intersect a single strip.
@@ 4.2.1 Existence of a Supersequence for Disjoint Unit Disks
Proof 4.16 (Proof of Lemma 4.15).
@@
    every vertex of the $Q$-hull is in distance 1 from the spine hull, or, in other words, every vertex of the $Q$-hull is contained in one of the strips, which follows from Observation 7.
@@
Proof 4.17 (Proof of Theorem 4.12).
@@
4.2.2 Computation of a Supersequence for Disjoint Unit Disks
In this section, we argue that we can also compute such a sequence $\Xi$ in $O(n\log n)$ time.
Theorem 4.18.
Let $\mathcal{R}$ be a set of $n$ disjoint unit disks in the plane. A $(1,3)$-smooth quarter-hull-supersequence $\Xi$ of $\mathcal{R}$ of size $O(n)$ can be computed in $O(n\log n)$ time.
First, we want to argue that we can compute the arrangement of all disks and strips in $O(n\log n)$ time.